OpenStudy (anonymous):

he graph of the function g(x)=2/3(x-3)^2(x+5)^4 has a zero of multiplicity 4 at x = -5 . True or False?

OpenStudy (anonymous):

do you know how to find zeros?

OpenStudy (valpey):

Is this what you are asking: \[g(x) = \frac{2}{3} (x-3)^2 + (x+5)^4\] has a zero of multiplicity 4 at x = -5?

OpenStudy (valpey):

Or the product? \[g(x)=\frac{2}{3}(x−3)^{2}(x+5)^{4}\]

OpenStudy (valpey):

I think the answer is true if the second one. because (x+5) = 0 and it is taken to the fourth power.